OpenStudy (anonymous):

Solve the equation. 2(x - 3)2 = 8 A) -1, -8 B) 1, 5 C) 7 D) 8

OpenStudy (anonymous):

open the brackets first! this time do it on ur own! gonna help if u are wrong or stuck! u can do it!

OpenStudy (anonymous):

im stuck ... for some reason im getting a way higher number

HanAkoSolo (jamierox4ev3r):

you can also use distribution

HanAkoSolo (jamierox4ev3r):

2(x-3) 2(x-3)

HanAkoSolo (jamierox4ev3r):

2x-6+2x-6=8

HanAkoSolo (jamierox4ev3r):

4x-12=8 +12 +12 4x=20 x=5

OpenStudy (anonymous):

@Jamierox4ev3r that's a quadratic equation ! u will get two answers which i guess are 1 and 5

HanAkoSolo (jamierox4ev3r):

right thats correct

OpenStudy (anonymous):

Thanks i see now

OpenStudy (anonymous):

so first if we open the bracket we get \[x ^{2}-6x+9=4\]